OpenStudy (anonymous):

Which is a difference of cubes? x6 – 27 x15 – 36 x16 – 64 x5 – 125

OpenStudy (anonymous):

@Nnesha

Nnesha (nnesha):

which one is a perfect cube root ???

Nnesha (nnesha):

which number**```

OpenStudy (anonymous):

125

Nnesha (nnesha):

there is one more

OpenStudy (anonymous):

36

Nnesha (nnesha):

what is \[\sqrt{36}\] and what is \[\sqrt[3]{36}\] which one give you perfect number ??

OpenStudy (anonymous):

6

Nnesha (nnesha):

for which give you 6 first one or second one ??

OpenStudy (anonymous):

first

OpenStudy (anonymous):

3 second

Nnesha (nnesha):

what ??

OpenStudy (anonymous):

3 Is The Second

Nnesha (nnesha):

square root of 36 give you 6 not the second one 36 IS NOT A PERFECT CUBE ROOT

OpenStudy (anonymous):

So D

Nnesha (nnesha):

in order to be a answer the degree of variable should also be a PERFECT CUBE ROOT

Nnesha (nnesha):

DO YOU HAVE CALCULATOR ??

OpenStudy (anonymous):

no

Nnesha (nnesha):

\[\sqrt[3]{ X ^{2}} ---> (X)^{2/3}\] change that x to x^5/3 can you get perfect number when you divide 5/3 ??